# Shadows and Observational Images of a Schwarzschild-like Black Hole Surrounded by a Dehnen-type Dark Matter Halo

### v1

@@ V The images of a Schwarzschild-like BH with static thin spherical accretions
Then, we investigated the optical appearance of a Schwarzschild-like BH illuminated by an optically thin accretion disk. Through three simplified models, we found that different models exhibit significant differences in the distribution of observed intensity and the resulting optical appearance. The light intensity contributions may arise from any combination of direct emission, lensing ring, and photon ring. Moreover, the DM halo parameters $\rho_{s}$ and $r_{s}$ significantly affect the distribution of bright rings in the optical image. As $\rho_{s}$ and $r_{s}$ increase, the associated characteristic radii also increase, leading to an expansion of the image radius and causing the bright rings to appear farther from the BH, thereby producing optical features that deviate from those of a Schwarzschild BH.
Finally, we further examined the optical characteristics of a Schwarzschild-like BH under a static spherical accretion scenario. The findings indicate that, akin to the case of an optically thin disk, the peak of the observed intensity moves to higher values of the impact parameter $b$ as $\rho_{s}$ and $r_{s}$ grow. However, the observed intensity gradually decreases, resulting in a diminished overall brightness of the optical image. These findings provide a valuable reference for understanding the variation patterns of the optical features of Schwarzschild-like BHs under the combined influence of specific accretion environments and DM halo parameters.

### v2

@@ 6 Conclusion
Then, we investigated the optical appearances of a Schwarzschild-like BH illuminated by optically thin accretion disks. Through three simplified models, we found that different models exhibit significant differences in the distribution of observed intensity and the resulting optical appearance, as shown in Figs. 9 and  10. The observed intensity profile can be decomposed into three distinct components: direct emission, lensing ring, and photon ring contributions. Moreover, the DM halo parameters $\rho_{s}$ and $r_{s}$ significantly affect the distribution of bright rings in the optical image. As $\rho_{s}$ and $r_{s}$ increase, the associated characteristic radius also increase, leading to an expansion of the image radius and causing the bright rings to appear farther from the BH, thereby producing optical features that deviate from those of a Schwarzschild BH.
Finally, we further examined the optical characteristics of a Schwarzschild-like BH under both static and radially infalling spherical accretion scenarios (see Figs. 12 and  14). The findings indicate that, akin to the case of an optically thin disk, the peak of the observed intensity moves to higher values of the impact parameter $b$ as $\rho_{s}$ and $r_{s}$ grow. However, the observed intensity gradually decreases, resulting in a diminished overall brightness of the optical image. These findings provide a valuable reference for understanding the variation patterns of the optical features of Schwarzschild-like BHs under the combined influence of specific accretion environments and DM halo parameters.
It is worth noting that extensive research has been conducted on BH solutions surrounded by various DM models Al-Badawi:2025njy ; Konoplya:2022hbl ; Konoplya:2025mvj ; Macedo:2024qky ; Li:2024abk ; Haroon:2025rzx ; Pantig:2022whj ; Liu:2023oab ; Liu:2022ygf . In studies concerning the shadows and optical appearance of BHs surrounded by DM Li:2024abk ; Zeng:2025kqw ; He:2024amh ; Jha:2025xjf ; Jha:2024ltc ; Al-Badawi:2024qpt ; Hou:2018avu ; Xu:2018mkl , the Schwarzschild BH embedded in Dehnen-(1, 4, 5/2) type DM—as well as its counterpart in the presence of a quintessence field Hamil:2025pte —exhibits a gradual increase in both shadow size and optical appearance as the DM parameter grows. This behavior aligns with shadow variations induced by certain DM models, such as cold DM Zeng:2021mok ; Zeng:2025kqw . In contrast, some DM models—for instance, Hernquist-type DM Macedo:2024qky —demonstrate a decrease in shadow size with increasing DM parameter, exhibiting trends distinct from those of the Dehnen-type model. These observations suggest that the optical appearance of a Schwarzschild-like BH illuminated under different accretion disk models may allow us to distinguish between various DM profiles based on image characteristics. For models exhibiting similar shadow behavior under parameter variation, further in-depth analysis will be required. A systematic comparison of optical appearances across different DM scenarios may be pursued in future research.